# Three Dimensional Topological Field Theories and Nahm Sum Formulas

### v1

@@ A.3 Modular data of A/B twisted TFTs
up to an overall phase factor. Alternatively, as a finite semisimple 3D TFT, the partition function can be given as
@@
with a $\delta\in\mathbb{Q}$ [^21] The phase factor $e^{i\pi\delta}$ comes from 3-manifold framing choice, background R-symmetry Chern-Simons term, gravitational Chern-Simons term and etc. All of these factors contribute to a rational value of $\delta$. We define
~$\begin{split}&\widetilde{{\cal H}}(\hat{u}_{\alpha},v):={\cal H}(\hat{u}_{\alpha},v)\prod_{i}\Pi^{-2v_{i}}_{i}(\hat{u}_{\alpha},v),\\ &\widetilde{{\cal F}}(\hat{u}_{\alpha},v):={\cal F}(\hat{u}_{\alpha},v)\prod_{i}\Pi^{v_{i}}_{i}(\hat{u}_{\alpha},v).\end{split}$ (105)
We expect an injective map from $[W_{\vec{Q}_{\alpha}}]\in W^{A}_{\rm sim}/\sim$, where $\sim$ is the IR equivalence in (36), to $\hat{u}_{\alpha}\in{\cal S}_{\rm BE}(v=-\frac{\mu_{0}-a}{2},\nu_{R}=-\frac{1}{2})$ satisfying
[…]
Using this relation, one can fix the matrix elements $S_{\alpha\beta}$ Cho:2020ljj ; Gang:2021hrd .
@@ A.4 Relation to Nahm’s conjecture
~$\begin{split}\widetilde{{\cal F}}(\hat{u},v)=e^{\frac{1}{2\pi i}\sum_{a}L(x_{a})}.\end{split}$ (109)
@@
|~$\begin{split}|\widetilde{{\cal F}}(\hat{u},v)|&=\exp\left(\frac{1}{2\pi}\sum_{a}{\rm Im}(\textrm{Li}_{2}(x_{a}))+\arg(1-x_{a})\log|x_{a}|\right)=e^{\frac{1}{2\pi}\sum_{a}D(x_{a})}.\end{split}$ (110)
@@ Appendix B Characters of rational VOAs
Virasoro minmal models:
[…]
@@
Super-Virasoro minimal models:
[…]

### v2

@@ A.3 Modular data of A/B twisted TFTs
up to an overall phase factor. As a finite semisimple 3d TFT, on the other hand, the partition function can be alternatively given as
@@
with a $\delta\in\mathbb{Q}$. [^22] The phase factor $e^{i\pi\delta}$ comes from 3-manifold framing choice, background R-symmetry Chern-Simons term, gravitational Chern-Simons term and etc. All of these factors contribute to a rational value of $\delta$. We define
$\begin{split}&\widetilde{{\cal H}}(\hat{u}_{\alpha},v):={\cal H}(\hat{u}_{\alpha},v)\prod_{i}\Pi^{-2v_{i}}_{i}(\hat{u}_{\alpha},v),\\ &\widetilde{{\cal F}}(\hat{u}_{\alpha},v):={\cal F}(\hat{u}_{\alpha},v)\prod_{i}\Pi^{v_{i}}_{i}(\hat{u}_{\alpha},v).\end{split}$ (114)
We expect an injective map from $W_{\vec{Q}_{\alpha}}]\in W^{A}_{\rm sim}/\sim$, where $\sim$ is the IR equivalence in (41), to $\hat{u}_{\alpha}\in{\cal S}_{\rm BE}(v=-\frac{\mu_{0}-a}{2},\nu_{R}=-\frac{1}{2})$ satisfying
$\prod\exp(-2\pi i\vec{Q}_{\alpha}\cdot\hat{u}_{\beta})=\left(\frac{S_{\alpha\beta}}{S_{0\beta}}\textrm{ of A-twisted TFT}\right)\;.$ (115)
Using the relation, one can fix the matrix element $S_{\alpha\beta}$ Cho:2020ljj ; Gang:2021hrd .
@@ A.4 Relation to Nahm’s conjecture
$\begin{split}\widetilde{{\cal F}}(\hat{u},v)=e^{\frac{1}{2\pi i}\sum_{a}L(x_{a})}.\end{split}$ (118)
@@
$\begin{split}|\widetilde{{\cal F}}(\hat{u},v)|&=\exp\left(\frac{1}{2\pi}\sum_{a}{\rm Im}(\textrm{Li}_{2}(x_{a}))+\arg(1-x_{a})\log|x_{a}|\right)=e^{\frac{1}{2\pi}\sum_{a}D(x_{a})}.\end{split}$ (119)
@@ Appendix B Characters of rational VOAs
The characters of Virasoro minimal model $M(P,Q)$ are ($1\leq r<P,\;1\leq s<Q$)
@@
The characters of ${\cal N}=1$ super Virasoro minimal model $SM(P,Q)$ are ($1\leq r<P,\;1\leq s<Q$ with $r-s\in 2\mathbb{Z}$)